# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Abstract
We constrain a broad class of teleportation protocols using insights from locality. In the “standard” teleportation protocols we consider, all outcome-dependent unitaries are Pauli operators conditioned on linear functions of the measurement outcomes. We find that all such protocols involve preparing a “resource state” exhibiting symmetry-protected topological (SPT) order with Abelian protecting symmetry $\mathcal{G}_{k}=(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{k}$. The $k$ logical states are teleported between the edges of the chain by measuring the corresponding $2k$ string order parameters in the bulk and applying outcome-dependent Paulis. Hence, this single class of nontrivial SPT states is both necessary and sufficient for the standard teleportation of $k$ qubits. We illustrate this result with several examples, including the cluster state, variants thereof, and a nonstabilizer hypergraph state.
@@ 1.1 Background
Entangled states of many-body quantum systems are valuable resources for a variety of tasks in quantum information processing. A prototypical example is quantum teleportation: By measuring an entangled state and applying unitaries conditioned on the outcomes (i.e., feedback), quantum information can be transferred over larger distances than can be achieved using unitary time evolution alone Bennett et al. (1993). Teleportation protocols can also be concatenated to form a quantum repeater that transfers quantum information over arbitrary distances Briegel et al. (1998).
It is then natural to ask what types of entangled many-body resource states can be used to teleport quantum information by measuring from one end of a 1D chain to the other and applying outcome-dependent feedback, as illustrated in Fig. 1. In fact, such entangled resource states form the backbone of many notions of quantum computation, and especially measurement-based quantum computation (MBQC), in which logical operations are applied to a quantum state as it is teleported Gottesman and Chuang (1999); Raussendorf and Briegel (2001). These same states are also closely related to the concept of localizable entanglement Popp et al. (2005), which has important connections to the ground-state properties and phase transitions of quantum spin chains Verstraete et al. (2004a, b); Jin and Korepin (2004); Pachos and Plenio (2004); Venuti and Roncaglia (2005); Skrøvseth and Bartlett (2009). Accordingly, significant effort has been invested into characterizing entangled quantum states by their utility as resource states for tasks such as MBQC and quantum teleportation Raussendorf et al. (2005); Van den Nest et al. (2006); Gross and Eisert (2007); Gross et al. (2009); Bremner et al. (2009); Doherty and Bartlett (2009); Gross and Eisert (2010); Else et al. (2012); Prakash and Wei (2015); Miller and Miyake (2016); Stephen et al. (2017); Marvian (2017); Wei (2018).
@@
Figure 1: Illustration of the central result of this paper: All standard teleportation protocols amount to preparing an SPT state with symmetry $\mathcal{G}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ with a logical state $\lvert\psi\rangle$ encoded (in purple), measuring the corresponding pair of string order parameters, and applying outcome-dependent Paulis. Teleporting $k\geq 1$ qubits requires measuring $2k$ strings.
Of particular interest is the class of many-body states in 1D with symmetry-protected topological (SPT) order Chen et al. (2011); Pollmann et al. (2010); Schuch et al. (2011); Chen et al. (2013), which are known to be useful resource states for quantum teleportation and MBQC Doherty and Bartlett (2009); Miyake (2010); Else et al. (2012); Prakash and Wei (2015); Stephen et al. (2017); Marvian (2017); Wei (2018). These states are short-range entangled, yet cannot be connected to product states via finite-depth unitaries that respect the “protecting” symmetry group $\mathcal{G}$. For an Abelian symmetry $\mathcal{G}$, the corresponding SPT states exhibit perfect long-range order characterized by nonlocal stringlike objects. These string order parameters act as commuting symmetry operators in the interior of any finite interval, but are decorated by endpoint operators at the boundaries den Nijs and Rommelse (1989); Kennedy and Tasaki (1992); Pollmann and Turner (2012). Moreover, measuring the symmetry charge in any such bulk region reduces the string order parameter to a long-range two-point correlation function of the two endpoint operators, leaving behind a pattern of long-range entanglement that can be used for teleportation Marvian (2017). Remarkably, a state’s suitability for teleportation is a property of entire SPT phases, and not merely special points within the phase Doherty and Bartlett (2009); Else et al. (2012); Miyake (2010); Raussendorf et al. (2022), establishing a compelling link between quantum information processing and quantum phases of matter.
Another source of insight into quantum information processing is locality, which imposes bounds on the speed with which quantum information can be transferred. The Lieb-Robinson Theorem Lieb and Robinson (1972) establishes an emergent speed limit in the context of unitary time evolution under generic local Hamiltonians. Additionally, more general quantum channels captured by local Lindbladians obey the same bound Poulin (2010). However, when local projective measurements are combined with nonlocal feedback (via instantaneous classical signalling), the standard Lieb-Robinson bound can be overcome. Nevertheless, a similar bound on generic useful quantum tasks has recently been derived even in the presence of nonlocal feedback Friedman et al. (2022a). These results impose optimal quantitative constraints on the distance over which logical qubits can be teleported.
@@ 1.2 Summary of assumptions and results
The main result of this work is that a class of “standard” teleportation protocols—which includes most, if not all, teleportation protocols known to the literature—is equivalent to preparing a state with a particular SPT order, measuring its string order parameters, and applying outcome-dependent Pauli gates to the destination qubits.
@@
We note that the overview in Sec. 2 is generic, in that it only assumes that the system of interest contains $N$ qubits and that all unitary operations act locally. The statements and proofs therein may also be found in the literature, and hold beyond the scope of this work. For example, we specify the intuitive notion of “physical” teleportation protocols in Def. 3, and briefly explain the bounds Lieb and Robinson (1972); Friedman et al. (2022a) that constrain them in Sec. 2.5.
@@
The main assumptions appear in Sec. 3. We formally define the “standard” teleportation of $k\geq 1$ logical qubits in Def. 6. We require that these protocols achieve a teleportation distance $L$ (13) that cannot be achieved using unitary evolution alone, to distinguish them from quantum state transfer (see Def. 1). We also restrict all measurements to single-qubit observables, as is standard in quantum information processing Nielsen and Chuang (2010). We assume that no physical unitaries are applied to previously measured qubits; such unitaries have no direct utility to teleportation, but needlessly complicate the statement of our main results. Our crucial assumption—which we expect is directly related to our main results—is that all quantum operations conditioned on the outcomes of prior measurements correspond to physical Pauli gates, and all classical side processing involves only linear functions of the measurement outcomes. We expect the extension of this analysis beyond these restrictions to lead to new classes of teleportation protocols with analogous results; we defer such an investigation to future work.
@@
In Sec. 4, we constrain standard teleportation protocols in several ways. We first show that at least two measurements—along with outcome-dependent feedback—are required to beat purely unitary protocols. We then show that two measurements are required per “measurement region,” with each measurement region no larger than $2v(T-1)$, where $v$ is the Lieb-Robinson velocity Lieb and Robinson (1972) and $T$ is the duration of time evolution or depth. Generic teleportation protocols involve concatenating measurement regions using pairs of measurements. Importantly, we show that all measured observables must commute, and provide bounds on teleportation with $M$ measurements. These results hold for any $k\geq 1$.
@@
In Sec. 5, we prove that the resource state $\lvert\Psi_{t}\rangle$ is an SPT with protecting symmetry $\mathbb{Z}_{2}^{2k}$. For $k=1$, this state is in the same phase as the cluster state Briegel and Raussendorf (2001). We also show that the measured observables define the bulk operators of $2k$ string order parameters that define the SPT order. Essentially, measuring the bulk of each pair of string order parameters leads to localizable entanglement that can be used to teleport a single logical qubit Popp et al. (2005); Marvian (2017). The general form of such protocols is depicted schematically in Fig. 1.
Importantly, while it was previously known that generic 1D SPT states could be used as resources for quantum teleportation, our results conversely show that all standard teleportation protocols imply a single class of SPT resource state. This establish an equivalence between (standard) teleportation and SPT states, a direct connection between a particular quantum phase of matter and a useful quantum tasks, and lends insight into the physical mechanism underlying quantum teleportation. In Sec. 6 we highlight these results using various examples, including nonstabilizer states and states for which the SPT order was not previously identified. Finally, in Sec. 7, we summarize our main results and comment on protocols beyond standard teleportation and future directions.
@@ 2 Teleportation overview
Here we briefly review the task of quantum teleportation. In Sec. 2.1 we identify the physical systems of interest, corresponding to qubits on a graph $G$, along with various useful operators. In Sec. 2.2, we define quantum state transfer—the measurement-free analogue of teleportation—and prove several results. In Sec. 2.3, we review the Stinespring representation Barberena and Friedman (2023); Friedman et al. (2022a, b); Stinespring (1955); Choi (1975); Takesaki (2012) of measurements and outcome-dependent operations as unitary channels on a dilated Hilbert space. In this sense, teleportation can be viewed as state transfer on the dilated Hilbert space, as we explain in Sec. 2.4, where we also consider an example of teleportation using the 1D cluster state Briegel and Raussendorf (2001), highlighting the Stinespring representation. Lastly, in Sec. 2.5, we present Theorem 2 due to Lieb and Robinson Lieb and Robinson (1972), which establishes a bound (28) on quantum state transfer. We then present Theorem 3 Friedman et al. (2022a), which extends Theorem 2 to generic quantum channels (involving measurements and nonlocal feedback) with a modified bound (29), and constrains teleportation.
@@ 2.1 Preliminaries
$\mathcal{H}_{\rm ph}=\bigotimes\limits_{v\in V}\,\mathbb{C}^{2}=\mathbb{C}^{2^{N}}\,,~{}~{}$ (1)
and dimension $\mathcal{D}_{\rm ph}=\operatorname*{dim}(\mathcal{H}_{\rm ph})=2^{N}$. We denote Pauli-string operators acting on $\mathcal{H}_{\rm ph}$ as, e.g.,
$\Gamma_{\boldsymbol{n}}=\bigotimes\limits_{v\in V}\,\sigma^{n_{v}}_{v}\equiv\sigma^{\boldsymbol{n}},~{}~{}$ (2)
where $\boldsymbol{n}=\{\,n_{v}\,|\,\forall\,v\in V\,\}$ is a length-$N$ vector that encodes the Pauli content $n_{v}\in\{0,1,2,3\}$ on each vertex $v\in V$, and $\sigma^{0}_{v}\equiv\mathds{1}$, $\sigma^{1}_{j}=\sigma^{x}_{j}=X_{j}$, and so on. The $4^{N}$ unique Pauli strings form a complete, orthonormal basis for operators acting on $\mathcal{H}_{\rm ph}$ (1); when multiplied by $\mathrm{i}^{k}$ (for $k=0,1,2,3$), these operators realize the Pauli group $\mathbf{P}_{*}{N}$. Relatedly, the Clifford group $\mathbf{C}_{*}{N}$ is the set of unitary operators that map the Pauli group $\mathbf{P}_{*}{N}$ to itself. Note that we frequently label both the Clifford and Pauli groups by the vertex set or Hilbert space to which they correspond (e.g., $\mathbf{C}_{*}{N}=\mathbf{C}_{*}{V}=\mathbf{C}_{*}{\rm ph}$).
Additionally, for a given $N$-qubit state $\lvert\Psi\rangle\in\mathcal{H}_{\rm ph}$, we define the unitary stabilizer group
@@
as the set of all unitaries $\mathsf{S}$ for which $\lvert\Psi\rangle$ is a $+1$ eigenstate (note that $\mathcal{S}$ always contains the identity $\mathds{1}$). Crucially, we do not require that $\mathcal{S}$ (3) be a subset of the Pauli group $\mathbf{P}_{*}{N}$—i.e., we do not restrict to “stabilizer states” $\lvert\Psi\rangle$ Gottesman (1997), for which all $\mathsf{S}\in\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$ are Pauli strings (2). In fact, in Sec. 6.2, we explicitly consider teleportation protocols involving a nonstabilizer hypergraph resource state Gottesman and Chuang (1999); Rossi et al. (2013); Gachechiladze et al. (2019); Miller and Miyake (2016). Moreover, we do not require that $\mathcal{S}$ (3) be Abelian (which is guaranteed when $\mathcal{S}\subset\mathbf{P}_{*}{N}$).

### v3

@@ Abstract
We constrain a broad class of teleportation protocols using insights from locality. In the “standard” teleportation protocols we consider, all outcome-dependent unitaries are Pauli operators conditioned on linear functions of the measurement outcomes. We find that all such protocols involve preparing a “resource state” exhibiting symmetry-protected topological (SPT) order with Abelian protecting symmetry $\mathcal{G}_{k}=(\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2})^{k}$. The $k$ logical states are teleported between the edges of the chain by measuring the corresponding $2k$ string order parameters in the bulk and applying outcome-dependent Paulis. Hence, this single class of nontrivial SPT states is both necessary and sufficient for the standard teleportation of $k$ qubits. We illustrate this result with several examples, including the cluster state, variants thereof, and a nonstabilizer hypergraph state.
@@ 1.1 Background
Entangled states of many-body quantum systems are valuable resources for a variety of tasks in quantum information processing. A prototypical example is quantum teleportation: By measuring an entangled state and applying unitaries conditioned on the outcomes (i.e., feedback), quantum information can be transferred over larger distances than can be achieved using unitary time evolution alone [1]. Teleportation protocols can also be concatenated to form a quantum repeater that transfers quantum information over arbitrary distances [2].
It is then natural to ask what types of entangled many-body resource states can be used to teleport quantum information by measuring from one end of a 1D chain to the other and applying outcome-dependent feedback, as illustrated in Fig. 1. In fact, such entangled resource states form the backbone of many notions of quantum computation, and especially measurement-based quantum computation (MBQC), in which logical operations are applied to a quantum state as it is teleported [3, 4]. These same states are also closely related to the concept of localizable entanglement [5], which has important connections to the ground-state properties and phase transitions of quantum spin chains [6, 7, 8, 9, 10, 11]. Moreover, the effect of imperfect teleportation on many-body ground states has recently been studied and implemented on quantum devices [12, 13, 14, 15]. Accordingly, significant effort has been invested into characterizing entangled quantum states by their utility as resource states for tasks such as MBQC and quantum teleportation [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@
Figure 1: Illustration of the central result of this paper: All standard teleportation protocols amount to preparing an SPT state with symmetry $\mathcal{G}=\mathbb{Z}^{\,}_{2}\times\mathbb{Z}^{\,}_{2}$ with a logical state $\lvert\psi\rangle$ encoded (in purple), measuring the corresponding pair of string order parameters, and applying outcome-dependent Paulis. Teleporting $k\geq 1$ qubits requires measuring $2k$ strings.
Of particular interest is the class of many-body states in 1D with symmetry-protected topological (SPT) order [31, 32, 33, 34], which are known to be useful resource states for quantum teleportation and MBQC [21, 22, 35, 25, 23, 24, 26, 27, 28]. These states are short-range entangled, yet cannot be connected to product states via finite-depth unitaries that respect the “protecting” symmetry group $\mathcal{G}$. For an Abelian symmetry $\mathcal{G}$, the corresponding SPT states exhibit perfect long-range order characterized by nonlocal stringlike objects. These string order parameters act as commuting symmetry operators in the interior of any finite interval, but are decorated by endpoint operators at the boundaries [36, 37, 38]. Moreover, measuring the symmetry charge in any such bulk region reduces the string order parameter to a long-range two-point correlation function of the two endpoint operators, leaving behind a pattern of long-range entanglement that can be used for teleportation [27]. Remarkably, a state’s suitability for teleportation is a property of entire SPT phases, and not merely special points within the phase [21, 23, 35, 39], establishing a compelling link between quantum information processing and quantum phases of matter.
Another source of insight into quantum information processing is locality, which imposes bounds on the speed with which quantum information can be transferred. The Lieb-Robinson Theorem [40] establishes an emergent speed limit in the context of unitary time evolution under generic local Hamiltonians. Additionally, more general quantum channels captured by local Lindbladians obey the same bound [41]. However, when local projective measurements are combined with nonlocal feedback (via instantaneous classical signalling), the standard Lieb-Robinson bound can be overcome. Nevertheless, a similar bound on generic useful quantum tasks has recently been derived even in the presence of nonlocal feedback [42]. These results impose optimal quantitative constraints on the distance over which logical qubits can be teleported.
@@ 1.2 Summary of assumptions and results
The main result of this work is that a standard class of teleportation protocols—which includes most, if not all, teleportation protocols known to the literature—is equivalent to preparing a state with a particular SPT order, measuring its string order parameters, and applying outcome-dependent Pauli gates to the destination qubits. We summarize our assumptions and main results below.
Summary of assumptions. We consider “standard” teleportation protocols $\mathcal{W}$ with the following properties:
  • The $k$ logical states $\lvert\psi_{n}\rangle$ are localized to single sites before and after the protocol $\mathcal{W}$.
  • Each qubit is teleported a distance greater than that which can be achieved using unitaries alone.
  • The unitary part of $\mathcal{W}$ is local in the sense of obeying a Lieb-Robinson bound [40].
  • The nonunitary part of $\mathcal{W}$ is limited to projective measurements and outcome-dependent “recovery” operations. We note that other operations cannot enhance teleportation [42].
  • All measured observables $A_{j}$ are unitarily connected to single-qubit observables.
  • No unitaries act on previously measured sites.
  • Finally, the most important assumptions are that: (i) the recovery operations are linear functions of the measurement outcomes and (ii) act on the physical qubits as Pauli operators or strings thereof.
These criteria are also summarized in Def. 6 of standard teleportation. The first five assumptions simply define the precise notion of teleportation we consider. The penultimate, technical criterion is imposed to simplify proofs, but it can almost certainly be relaxed. Most importantly, the final criterion is our main conceptual assumption, and we find that it strongly constrains the class of teleportation protocols and corresponding resource states. We expect that relaxing this final assumption in various ways will lead to different classes of SPT resource states beyond what we explore herein.
For convenience, we write any such protocol $\mathcal{W}$ in the “canonical form” $\mathcal{W}=\mathcal{R}\,\mathcal{M}\,\mathcal{U}$ (31), where $\mathcal{U}$ acts unitarily (see Sec. 3.1). We view $\mathcal{U}$ as preparing a “resource” state (36), which is then subjected to the nonunitary measurement and recovery channels (respectively, $\mathcal{M}$ and $\mathcal{R}$) to achieve standard teleportation.
Summary of results. Our main results for such standard teleportation protocols are as follows:
  • The distance $L$ (13) over which $k$ qubits can be teleported using $M$ measurements is bounded by
    $L\leq vT+2\,v\left\lfloor\frac{M}{2k}\right\rfloor\left(T-1\right)\,,~{}$ (47)
    with the additional requirement $T\geq 1+k/v$ (48), as established in Theorem 13.
  • All measured observables in the canonical-form protocol must commute with one another and do not act on the final logical sites, as proven in Cor. 14.
  • If $\mathcal{W}$ teleports $k$ qubits as described above acting on an initial state $\lvert\Psi_{0}\rangle$, then the resource state $\lvert\Psi_{t}\rangle=\mathcal{U}\lvert\Psi_{0}\rangle$ (36) has SPT order protected by a $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{k}$ symmetry. For $k=1$, e.g., all such states are unitarily equivalent to the cluster state [43]. This result appears in Theorem 18.
  • The resource state $\lvert\Psi_{t}\rangle$ (36) has perfect string order, and thus lies at the fixed point of the $(\mathbb{Z}_{2}\times\mathbb{Z}_{2})^{k}$ SPT phase [44]. Each of the $2k$ string order parameters is the product of all measured observables used to determine the $X$-type (or $Z$-type) recovery operation for each of the $k$ logical qubits.
  • The resource state need not be a stabilizer state, as with the hypergraph state we consider in Sec. 6.2.
Importantly, while it was previously known that generic 1D SPT states could be used as resources for quantum teleportation, our results conversely show that all standard teleportation protocols imply a single class of SPT resource state. This establishes an equivalence between (standard) teleportation and SPT states, a direct connection between a particular quantum phase of matter and a useful quantum tasks, and lends insight into the physical mechanism underlying quantum teleportation.
1.3 Organization of the paper
@@
We begin with a generic overview in Sec. 2, in which we assume only that the system of interest contains $N$ qubits and that all unitary operations act locally. The statements and proofs therein may also be found in the literature, and hold beyond the scope of this work. For example, we specify the intuitive notion of “physical” teleportation protocols in Def. 3, and briefly explain the bounds [40, 42] that constrain them in Sec. 2.5.
@@
The main assumptions appear in Sec. 3. We formally define the “standard” teleportation of $k\geq 1$ logical qubits in Def. 6. We require that these protocols achieve a teleportation distance $L$ (13) that cannot be achieved using unitary evolution alone, to distinguish them from quantum state transfer (see Def. 1). We also restrict all measurements to single-qubit observables, as is standard in quantum information processing [45]. We assume that no physical unitaries are applied to previously measured qubits; such unitaries have no direct utility to teleportation, but needlessly complicate the statement of our main results. Our crucial assumption—which we expect is directly related to our main results—is that all quantum operations conditioned on the outcomes of prior measurements correspond to physical Pauli gates, and all classical side processing involves only linear functions of the measurement outcomes. We expect the extension of this analysis beyond these restrictions to lead to new classes of teleportation protocols with analogous results; we defer such an investigation to future work.
@@
In Sec. 4, we constrain standard teleportation protocols in several ways. We first show that at least two measurements—along with outcome-dependent feedback—are required to beat purely unitary protocols. We then show that two measurements are required per “measurement region,” with each measurement region no larger than $2v(T-1)$, where $v$ is the Lieb-Robinson velocity [40] and $T$ is the duration of time evolution or depth. Generic teleportation protocols involve concatenating measurement regions using pairs of measurements. Importantly, we show that all measured observables must commute, and provide bounds on teleportation with $M$ measurements. These results hold for any $k\geq 1$.
@@
In Sec. 5, we prove that the resource state $\lvert\Psi_{t}\rangle$ is an SPT with protecting symmetry $\mathbb{Z}_{2}^{2k}$. For $k=1$, this state is in the same phase as the cluster state [43]. We also show that the measured observables define the bulk operators of $2k$ string order parameters that define the SPT order. Essentially, measuring the bulk of each pair of string order parameters leads to localizable entanglement that can be used to teleport a single logical qubit [5, 27]. The general form of such protocols is depicted schematically in Fig. 1.
In Sec. 6 we highlight these results using various examples, including nonstabilizer states and states for which the SPT order was not previously identified. Finally, in Sec. 7, we summarize our main results, comment on protocols beyond standard teleportation, and discuss future directions and extensions of this work.
@@ 2 Teleportation overview
Here we briefly review the task of quantum teleportation. In Sec. 2.1 we identify the physical systems of interest, corresponding to qubits on a graph $G$, along with various useful operators. In Sec. 2.2, we define quantum state transfer—the measurement-free analogue of teleportation—and prove several results. In Sec. 2.3, we review the Stinespring representation [46, 42, 47, 48, 49, 50] of measurements and outcome-dependent operations as unitary channels on a dilated Hilbert space. In this sense, teleportation can be viewed as state transfer on the dilated Hilbert space, as we explain in Sec. 2.4, where we also consider an example of teleportation using the 1D cluster state [43], highlighting the Stinespring representation. Lastly, in Sec. 2.5, we present Theorem 2 due to Lieb and Robinson [40], which establishes a bound (28) on quantum state transfer. We then present Theorem 3 [42], which extends Theorem 2 to generic quantum channels (involving measurements and nonlocal feedback) with a modified bound (29), and constrains teleportation.
@@ 2.1 Preliminaries
$\mathcal{H}^{\vphantom{\prime}}_{\rm ph}=\bigotimes\limits_{v\in V}\,\mathbb{C}^{2}=\mathbb{C}^{2^{N}}\,,~{}~{}$ (1)
and dimension $\mathcal{D}^{\,}_{\rm ph}=\operatorname*{dim}(\mathcal{H}^{\,}_{\rm ph})=2^{N}$. We denote Pauli-string operators acting on $\mathcal{H}^{\,}_{\rm ph}$ as, e.g.,
$\Gamma^{\vphantom{\prime}}_{\boldsymbol{n}}=\bigotimes\limits_{v\in V}\,\sigma^{n^{\,}_{v}}_{v}\equiv\sigma^{\boldsymbol{n}}_{\,},~{}~{}$ (2)
where $\boldsymbol{n}=\{\,n^{\,}_{v}\,|\,\forall\,v\in V\,\}$ is a length-$N$ vector that encodes the Pauli content $n^{\,}_{v}\in\{0,1,2,3\}$ on each vertex $v\in V$, and $\sigma^{0}_{v}\equiv\mathds{1}$, $\sigma^{1}_{j}=\sigma^{x}_{j}=X_{j}$, and so on. The $4^{N}$ unique Pauli strings form a complete, orthonormal basis for operators acting on $\mathcal{H}^{\,}_{\rm ph}$ (1); when multiplied by $\mathrm{i}^{k}$ (for $k=0,1,2,3$), these operators realize the Pauli group $\mathbf{P}^{\vphantom{*}}_{*}{N}$. Relatedly, the Clifford group $\mathbf{C}^{\vphantom{*}}_{*}{N}$ is the set of unitary operators that map the Pauli group $\mathbf{P}^{\vphantom{*}}_{*}{N}$ to itself. Note that we frequently label both the Clifford and Pauli groups by the vertex set or Hilbert space to which they correspond (e.g., $\mathbf{C}^{\vphantom{*}}_{*}{N}=\mathbf{C}^{\vphantom{*}}_{*}{V}=\mathbf{C}^{\vphantom{*}}_{*}{\rm ph}$).
Additionally, for a given $N$-qubit state $\lvert\Psi\rangle\in\mathcal{H}^{\,}_{\rm ph}$, we define the unitary stabilizer group
@@
as the set of all unitaries $\mathsf{S}$ for which $\lvert\Psi\rangle$ is a $+1$ eigenstate (note that $\mathcal{S}$ always contains the identity $\mathds{1}$). Crucially, we do not require that $\mathcal{S}$ (3) be a subset of the Pauli group $\mathbf{P}^{\vphantom{*}}_{*}{N}$—i.e., we do not restrict to “stabilizer states” $\lvert\Psi\rangle$ [51], for which all $\mathsf{S}\in\operatorname*{stab}\left(*\right){\lvert\Psi\rangle}$ are Pauli strings (2). In fact, in Sec. 6.2, we explicitly consider teleportation protocols involving a nonstabilizer hypergraph resource state [3, 52, 53, 25]. Moreover, we do not require that $\mathcal{S}$ (3) be Abelian (which is guaranteed when $\mathcal{S}\subset\mathbf{P}^{\vphantom{*}}_{*}{N}$).